# PRISM: Robust VLM Alignment with Principled Reasoning for Integrated Safety in Multimodality

### v1

@@ 4.3 Evaluation
A fundamental challenge in VLM safety is achieving the delicate balance between harmlessness and helpfulness. Models that are overly defensive may reject legitimate queries, diminishing their practical utility, while those prioritizing helpfulness risk responding to harmful requests. Our evaluation specifically addresses this dual objective by assessing both safety robustness against malicious attacks and performance preservation on benign tasks.
@@
Safety Evaluations. To perform a comprehensive safety evaluation, we assess the model across three established benchmarks, each designed to probe distinct multimodal safety vulnerabilities. The first, JailBreakV-28K Luo et al. (2024) is a pioneering benchmark that evaluates both language model-transferred attacks (LLM-Trans) and direct multimodal attacks (MLLM). The second is the Challenge subset of VLBreakBench Wang et al. (2024c), which utilizes an adaptive jailbreaking framework where Gemini-1.5-Pro iteratively generates transferable multimodal attacks against a GPT-4o-mini target, with adversarial images synthesized by Stable Diffusion 3 Esser et al. (2024). The third, the Multi-Image Safety (MIS) benchmark Ding et al. (2025b), introduces a significant challenge by requiring models to reason across multiple images to detect safety violations—a scenario that falls outside the alignment scope of alignment data. We report the Attack Success Rate (ASR), and it was evaluated by llama-guard-3-8B Inan et al. (2023) as the judge for JailbreakV and VLBreak and GPT-4o for MIS by following their settings.
Beyond these static benchmarks, adaptive attacks represent a critical yet often neglected dimension of safety evaluation, as they more accurately reflect real-world scenarios where attackers iteratively refine their strategies based on model responses. Unlike static benchmarks that employ fixed attack patterns, dynamic evaluation frameworks assess model robustness against evolving attack strategies that adaptively exploit model vulnerabilities. To evaluate model robustness against adaptive threats, we employ IDEATOR Wang et al. (2024c) as our dynamic attack framework, which iteratively optimizes adversarial prompts and images. Following the original implementation protocol, we utilize MiniGPT-4 as the attacking agent and evaluate on a harmful corpus: AdvBench Zou et al. (2023). We configure the search parameters with width $N_{w}=7$ and depth $N_{d}=3$, thus having a total of $21$ Attack Queries.
Baselines. We compare our PRISM framework against several key baseline methods for VLM safety. These include SafeRLHF-V Ji et al. (2025), which applies Direct Preference Optimization (DPO) on a dual preference dataset labeled for both helpfulness and safety. Similarly, SPA-VL Zhang et al. (2025b) also leverages DPO, but on a large-scale preference dataset collected from a wide range of models. We also include VLGuard Zong et al. (2024), a method based on Supervised Fine-Tuning (SFT) that uses a dedicated vision-language safety instruction-following dataset.
Helpfulness Evaluation. One critical aspect of safeguarding VLMs is balancing the harmlessness and helpfulness, ensuring that models can effectively assist users while maintaining robust defenses against harmful content. To evaluate this balance, we employ the MM-Vet-v2 benchmark Yu et al. (2024), which provides a comprehensive assessment of multimodal models across various tasks including reasoning, generation, OCR, spatial understanding, knowledge, sequential reasoning, and mathematical capabilities. The benchmark is scored by GPT-4 (gpt-4-0613) with a total score ranging from 0 to 100.
@@
Main Results. Our experimental results in Table 1 demonstrate that PRISM achieves a better balance between safety robustness and helpfulness preservation, successfully navigating the critical trade-off in VLM alignment. Across established safety benchmarks, PRISM consistently delivers top-tier performance. On the Qwen2-VL model, for example, it achieves the lowest ASR of $1.46\%$ on JailbreakV-28K (LLM-Trans) and a near-zero ASR on the VLBreak Challenge. While a method like VLGuard also achieves strong safety metrics, it comes at a significant cost to the model’s core utility. This is evident in VLGuard’s drastic drop in helpfulness, where its MM-Vet-v2 score plummets to just $17.7$ for Qwen2-VL, indicating severe over-defense and a loss of general capabilities.
In contrast, PRISM avoids this pitfall. It secures its robust safety profile while maintaining a high level of helpfulness, scoring $20.4$ and $48.9$ on MM-Vet-v2 for LLaVA-1.5 and Qwen2-VL respectively.
Moreover, as illustrated in Figure 5(a), our proposed method demonstrates significant robustness against adaptive attacks. The aligned model consistently maintains a low ASR and requires a large number of queries to find an adversarial example, which translates to a substantially higher computational cost for the attacker. In contrast, the undefended base model is far more vulnerable. Its ASR shows three apparent increases, which correspond to the attack’s architecture: with a search depth of $N_{d}=3$, the attacking agent analyzes past failures to refine its strategy after every $N_{w}=7$ iterations, causing these sharp jumps in ASR.
@@ 4.4 Test Time Scaling
The structured nature of the reasoning process enables performance enhancement through test-time scaling. We achieve this by using a self-rewarding method, wherein the model itself functions as a reward model to guide a step-wise scaling process. We then employ Beam Search and Best-of-N strategies to generate and select optimal reasoning paths Snell et al. (2025). The associated computational budget is estimated by the number of candidate comparisons at each reasoning step relative to a baseline without such scaling. By filtering those questions that successfully jailbreak more than half of the defense method in MIS, we get a MIS-Challenge subset to evaluate our PRISM-DPO’s scaling performance on Qwen2-VL.
@@
5 Analysis and Ablation
To gain a better understanding of the reason why our PRISM methods are effective, we conduct experiments of analysis and ablation studies.
@@ 5 Analysis and Ablation
LLaVA-1.5 ND 78.81 13.1
PRISM 20.00 20.4
LLaVA-1.6 ND 62.97 20.9
PRISM 16.02 27.9
@@
Our method’s generalizability was tested by extending to more VLMs, including LLaVA-1.6-7B and Qwen2.5-VL-7B. We follow the same training and evaluation settings as described in the previous sections. The results, detailed in Table 2(a), reveal that our PRISM framework exhibits remarkable and consistent effectiveness across different models. Alongside enhancing safety on the most challenging MIS-Hard benchmark with more than 75% ASR decrease, PRISM compromises minimally or even improves utility.
Moreover, to understand the importance of each part in our PRISM framework, we ran an ablation study on the PRISM-CoT dataset. For this test, we trained different versions of our PRISM-SFT model, each time removing one of the three data categories: Problem Unsafe (PU), Image Unsafe (IU), or Problem-Image Combination Unsafe (CU). The results, shown in Table 2(b), clearly indicate that all three parts are necessary to achieve a more comprehensive safety. The most serious drop in safety occurred when we removed the Combination Unsafe (CU) data. This single change caused the Attack Success Rate (ASR) on the MIS-Hard benchmark to jump to $55.25\%$. This result is especially important because the MIS-Hard benchmark is built to test for harmful intent that arises from the connection between the image and the text. Removing the PU and IU data also made the model less safe, confirming that all three data types are critical. This shows they work together to guard the model against many different types of misuse.
@@ 6 Conclusion
In this work, we introduce PRISM, a framework designed to protect Vision-Language Models from sophisticated multimodal attacks by safety-aware reasoning. By leveraging such a reasoning process that is refined with direct preference optimization, PRISM achieves state-of-the-art safety across a wide range of benchmarks, including challenging out-of-distribution scenarios. Crucially, these significant safety improvements are realized without degrading the model’s core utility, effectively resolving the critical trade-off between harmlessness and helpfulness. Our findings validate that embedding explicit reasoning is a highly effective and scalable strategy for building VLMs that are both robustly safe and genuinely capable.
@@ Appendix B Dataset Curation Details
In Figure 7, 12, 13, we provide the detailed prompt used in prompting the GPT-4o to generate corresponding reasoning stages with the given labels.
We provide the detailed prompts used to instruct GPT-4o to score the model-generated reasoning steps during the MCTS search. For the helpfulness evaluation, each step is scored against ground truth results from the original dataset using the prompt in Figure 14. The prompts for evaluating the safety score are given in Figures 15 and 16.
Appendix C Helpfulness Evaluation
@@
Appendix D Failure Case Analysis

### v2

@@ 4.1 Evaluation
Safety Benchmarks. We evaluate on three benchmarks measuring distinct safety dimensions: (1) JailbreakV-28K Luo et al. (2024), covering both LLM-transferred and direct multimodal attacks; (2) VLBreakBench Challenge Wang et al. (2024c), where Gemini-1.5-Pro iteratively generates transferable attacks with adversarial images from Stable Diffusion 3 Esser et al. (2024); and (3) MIS Ding et al. (2025), requiring multi-image reasoning to detect safety violations. ASR is evaluated JailbreakV, VLBreak and MIS, following their original settings. For adaptive attack evaluation, we employ IDEATOR Wang et al. (2024c) with MiniGPT-4 as the attacker on AdvBench Zou et al. (2023).
Baselines. We compare our proposed PRISM, with training details provided in Appendix E, against SafeRLHF-V Ji et al. (2025), which employs Direct Preference Optimization on dual-preference data. Additionally, we evaluate against SPA-VL Zhang et al. (2025b), which utilizes Direct Preference Optimization on preference data. Finally, we include VLGuard Zong et al. (2024), which applies Supervised Fine-Tuning on safety instruction data.
Helpfulness Evaluation. We use MM-Vet-v2 Yu et al. (2024) following their setting to assess utility preservation across reasoning, OCR, spatial understanding, and mathematics tasks.
@@
4.2 Main Results.
Table 1 shows that PRISM achieves a favorable trade-off between safety robustness and helpfulness retention. On Qwen2-VL, PRISM reduces JailbreakV-28K (LLM-Trans) ASR to $1.46\%$ and achieves $0.05\%$ ASR on the VLBreak Challenge, while maintaining substantially higher MM-Vet-v2 helpfulness than other safety-aligned baselines (e.g., $48.9$ vs. $46.8$ for SPA-VL and far above $17.7/12.9$ for VLGuard/SafeRLHF-V). On LLaVA-1.5, PRISM achieves consistently low ASR across JailbreakV-28K, VLBreak, and MIS, and attains the highest MM-Vet-v2 score among the evaluated defenses ($20.4$). Overall, while safety alignment can reduce absolute helpfulness for strong base models, PRISM substantially mitigates the utility loss commonly induced by aggressive refusal training, yielding robust safety with competitive benign performance.
Moreover, as illustrated in Figure 5(a), our proposed method demonstrates significant robustness against adaptive attacks. Our models consistently maintain a low ASR and require a large number of queries to find a successful adversarial example, which translates to a substantially higher computational cost for the attacker. In contrast, the undefended base model is far more vulnerable. Its ASR shows three apparent increases, which correspond to the attack’s architecture: with a search depth of $N_{d}=3$, the attacking agent analyzes past failures to refine its strategy after every $N_{w}=7$ iterations, causing these sharp jumps in ASR. However, our method can consistently maintain low ASR.
@@ 4.3 Test Time Scaling
The structured nature of the reasoning process enables performance enhancement through test-time scaling. We achieve this by using a self-rewarding method, wherein the model itself functions as a reward model to guide a step-wise scaling process. We then employ Beam Search and Best-of-N strategies to generate and select best reasoning paths Snell et al. (2025). The associated computational budget is estimated by the number of candidate comparisons at each reasoning step relative to a baseline without such scaling. By filtering those questions that successfully jailbreak more than half of the defense method in MIS, we get a MIS-Challenge subset to evaluate our PRISM-DPO’s scaling performance on Qwen2-VL.
@@
4.4 Analysis and Ablation
@@
Qwen3-VL ND 12.78 66.0
PRISM 2.68 62.5
@@
Effectiveness across different base models. We evaluate PRISM on multiple VLM backbones, using the same training protocol and hyperparameters as in prior experiments. As shown in Table 2(a), PRISM consistently reduces the MIS-Hard attack success rate (ASR) while largely preserving helpfulness on MM-Vet2. For example, on LLaVA-1.6, PRISM lowers MIS-Hard ASR from $62.97\%$ to $16.02\%$ and even improves MM-Vet2 from $20.9$ to $27.9$. On Qwen2.5-VL, ASR drops sharply from $64.16\%$ to $6.34\%$ with a moderate MM-Vet2 decrease. Notably, this trend also holds for more recent and already better-aligned backbones: Qwen3-VL starts with a much lower MIS-Hard ASR ($12.78\%$), yet PRISM still achieves a further $4.8\times$ reduction to $2.68\%$ while maintaining high MM-Vet2. Overall, these results suggest PRISM provides transferable safety gains beyond a single base model, including settings where the backbone is strong and comparatively well aligned.
Effectiveness of reasoning step designs. Moreover, to understand the importance of each part in our PRISM framework, we conduct an ablation study on the PRISM-CoT dataset. We trained different versions of our PRISM-SFT model, each time removing one of the three data categories: Problem Unsafe (PU), Image Unsafe (IU), or Problem-Image Combination Unsafe (CU). The results, shown in Table 2(b), clearly indicate that all three parts are necessary to achieve better safety. The most serious drop in safety occurred when we removed the Combination Unsafe (CU) data. This single change caused the Attack Success Rate (ASR) on the MIS-Hard benchmark to jump to $55.25\%$. Removing the PU and IU data also made the model less safe, confirming that all three data types are critical. This shows they work together to align the model against many different types of threats. We further ablate the contribution of each reasoning stage at inference time by appending end tokens to skip specific steps. Results in Appendix Table 4 confirm that every stage contributes: removing Reasoning causes the largest utility drop ($38.21$ vs. $41.27$ on MMMU-Pro), while removing both Problem and Caption causes the largest safety degradation .
Ablating MCTS Structure. Table 3 progressively ablates the key design choices in PRISM-DPO. SFT alone achieves only modest safety gains. Adding naive per-step DPO reduces ASR to $7.8\%$$6.9\%$ but severely degrades utility to $33.0$, falling into the over-defense trap. Introducing MCTS tree structure with safety back-propagation (w SBP) recovers utility to $43.1$, but safety improvement plateaus at $5.1\%$$11.0\%$ ASR—because propagating safety penalties to preceding steps creates incorrect credit assignment: a stage that reasons correctly should not be penalized for a later stage’s safety failure. PRISM-DPO resolves this by keeping safety rewards stage-local (no safety BP) while back-propagating helpfulness rewards, achieving the best of both: lowest ASR with highest utility. This confirms that stage-local safety reward assignment is the critical design principle behind PRISM-DPO’s superior safety–utility balance.
Training JBV-mini MIS $M$-Vet2
Method ASR $\downarrow$ ASR $\downarrow$ Acc $\uparrow$
SFT only 13.2 30.1 38.5
SFT + naive DPO 7.8 6.9 33.0
SFT + MCTS w SBP 5.1 11.0 43.1
SFT + PRISM-DPO 0.38 6.0 48.9
Table 3: Comparison of training configurations on Qwen2-VL. “MCTS w SBP” uses MCTS tree structure with safety reward back-propagation, propagating safety penalties to preceding reasoning steps. PRISM-DPO instead assigns safety rewards stage-locally (no safety BP) and back-propagates only helpfulness rewards, achieving the best safety–utility balance.
@@ 5 Conclusion
In this work, we introduce PRISM, a reasoning-based framework designed to protect VLMs from multimodal jailbreak attacks. Across a diverse set of safety benchmarks, PRISM consistently achieves strong robustness and often matches or surpasses prior training-based defenses, including under adaptive and out-of-distribution settings. At the same time, PRISM retains substantially higher benign utility than existing safety alignment baselines, indicating that structured multimodal reasoning can improve the safety–utility trade-off. Our findings suggest that embedding explicit, stage-structured reasoning is an effective and scalable direction for building VLMs that are both safer and practically useful.
Appendix
@@ Appendix B Dataset Curation Details
Appendix C Detailed PRISM-CoT Reasoning Step Descriptions
We employ GPT-4o to generate structured chain-of-thought reasoning processes through the following four-step framework:
PROBLEM. This initial step entails a comprehensive examination of the textual prompt to identify potential harmful content or malicious intent. We provide only the text prompt and instruct the model to analyze the underlying intention behind the query and determine whether any safety violations are present. Upon detection of a violation, the specific category of safety infringement is explicitly identified according to our established taxonomy.
CAPTION. Recognizing that effective image safety evaluation requires contextual understanding, the model generates a caption that describes the visual content specifically in relation to the problem context. This contextualized captioning approach facilitates the identification of safety violations that might be apparent only when considering the image within the framework of the accompanying query.
REASONING. This critical step involves the integration and synthesis of information derived from both the problem analysis and visual assessment. The model conducts a systematic evaluation of the interaction between textual and visual modalities, identifying potential safety concerns that emerge specifically from their combination, which is a capability essential for detecting sophisticated multimodal safety violations.
OUTPUT. Upon detection of any safety violation, an appropriate refusal response is generated that explicitly articulates the specific type of violation identified in the preceding analytical steps. This ensures alignment with established safety protocols while maintaining transparency regarding the reasoning behind the refusal.
This structured reasoning framework ensures coverage of potential safety violations across individual modalities and their interactions. When malicious intent is detected, subsequent reasoning steps maintain the structural integrity of the process while appropriately refusing to engage with harmful content, thereby creating valuable training examples for our fine-tuning procedure.
Curating Prompts. In Figure 7, 12, 13, we provide the detailed prompts used to generate the corresponding reasoning stages with the given labels.
We provide the detailed prompts used to score the model-generated reasoning steps during the MCTS search. For the helpfulness evaluation, each step is scored against ground truth results from the original dataset using the prompt in Figure 14. The prompts for evaluating the safety score are given in Figures 15 and 16.
Appendix D Helpfulness Evaluation
@@
Appendix E PRISM Training Details
As illustrated in Figure 3, we first fine-tune our models on the PRISM-CoT dataset for $5$ epochs, with each reasoning step delimited by specialized tokens. We augment the model vocabulary with eight special tokens to demarcate the structured reasoning process. The fine-tuning procedure involves comprehensive parameter optimization using $90\%$ of the dataset for training and $10\%$ for validation. We conduct training with a learning rate of $1\times 10^{-5}$. Following the SFT phase, we proceed with direct preference optimization (DPO) training on the PRISM-DPO dataset. This phase involves optimizing the model to achieve finer-grained alignment at step-level. For DPO implementation, we employ LoRA fine-tuning with rank $r=16$ and scaling factor $\alpha=64$. We utilize a batch size of $16$ and a learning rate of $1\times 10^{-5}$, conducting training for $3$ epochs.
Appendix F MCTS Implementation Details
We initiate the MCTS process with a curated dataset comprising 500 benign and 1,500 malicious image-text pairs. Our implementation employs $k=3$ candidate children for each node during the expansion phase, with a maximum iteration limit of $200$ per reasoning tree to ensure computational tractability while maintaining sufficient exploration. For preference pair generation, we establish a difference margin of $\epsilon=0.4$ and a quality threshold of $\theta=0.8$ to identify statistically significant reward differences between reasoning paths.
Appendix G Ablation on Structured Reasoning Stages
Setting MIS-hard ASR $\downarrow$ MMMU-Pro Acc $\uparrow$
Full PRISM-DPO 11.3 41.27
w/o PROBLEM 19.3 39.42
w/o CAPTION 14.5 40.00
w/o PROBLEM + CAPTION 29.5 36.07
w/o REASONING 20.0 38.21
Table 4: Ablation on structured reasoning stages. Removing any stage degrades both safety and utility, confirming each stage’s contribution.
To isolate whether safety gains stem from the structured reasoning process itself rather than additional supervision, we perform controlled ablations by perturbing each of the four stages (Problem, Caption, Reasoning, Output) via appending end tokens to skip specific steps. As shown in Table 4, the full PRISM pipeline achieves the best balance. Notably, the Reasoning stage is crucial; removing it causes a significant drop in benign utility ($38.21$ vs $41.27$), confirming that improvements reflect genuine reasoning ability rather than just data curation. Ablating any stage increases ASR on MIS-hard, and this is most notable when both Problem and Caption are removed ($29.5$).
Appendix H Comparison with STAIR
Models MMMU-Pro MIS-hard VLBreak
Acc $\uparrow$ ASR $\downarrow$ ASR $\downarrow$
Qwen2-VL (Base) 44.2 78.02 15.12
+ STAIR-1k 41.4 33.19 3.19
+ PRISM-1k 42.9 13.00 0.88
Table 5: Head-to-head comparison between PRISM and STAIR Zhang et al. (2025a) trained on identical 1k data. PRISM’s multimodal-specific design yields substantially lower ASR on cross-modal benchmarks, while also better preserving utility on MMMU-Pro.
STAIR Zhang et al. (2025a) is the closest related work in the text-only safety reasoning literature, following a two-stage SFT+DPO paradigm with MCTS-based preference data generation. While PRISM introduces three principled and complementary contributions that are specifically motivated by the unique challenges of multimodal safety alignment, which STAIR is not designed to address.
Multimodal-specific structured reasoning. STAIR operates on a linear chain-of-thought over textual input only. PRISM instead structures reasoning into four dedicated stages: Problem (analyzing textual intent), Caption (providing a query-aware visual description), Reasoning (cross-modal synthesis to detect interaction-level threats), and Output (generating a justified response). This is a key innovation absent from STAIR: it forces the model to ground its visual understanding relative to the query, which is essential for detecting combination-unsafe scenarios where malicious intent only emerges from the interplay between text and image.
Three-category safety coverage. STAIR’s training data targets text-based safety violations. PRISM explicitly covers three distinct threat categories: Problem-unsafe (harmful text), Image-unsafe (harmful visuals), and Combination-unsafe (benign text and image that become harmful together). This categorical decomposition enables the model to develop targeted reasoning strategies for each threat type, particularly for the most challenging combination-unsafe threats that require sophisticated cross-modal reasoning.
Stage-local safety reward assignment. A critical design choice in PRISM-DPO is that safety rewards are not back-propagated through the MCTS tree. In STAIR, rewards propagate globally across all reasoning steps. This is appropriate for text-only helpfulness objectives, but problematic for safety: a preceding reasoning step that correctly analyzes the context should not be penalized for a later step’s safety failure, as this constitutes incorrect credit assignment. PRISM instead assigns safety rewards locally at each stage, while back-propagating only helpfulness rewards. As demonstrated in Table 3, this design yields the best safety–utility balance.
We conduct a head-to-head comparison by faithfully re-implementing the vanilla STAIR pipeline and training it on the identical 1k data subset used for PRISM. As shown in Table 5, while both methods achieve comparable utility on MMMU-Pro ($41.4$ vs. $42.9$), the safety gap is substantial. STAIR-1k reduces MIS-hard ASR to $33.19\%$, still far above PRISM-1k’s $13.00\%$, a relative improvement of $2.6\times$. On VLBreak, PRISM-1k achieves $0.88\%$ ASR versus $3.19\%$ for STAIR-1k, a $3.6\times$ improvement. The degradation of STAIR on MIS-hard is particularly instructive: this benchmark tests precisely the combination-unsafe threat model where neither text nor image alone reveals the harmful intent. Without a dedicated visual grounding stage, cross-modal threat category coverage, or stage-local reward design, STAIR lacks the mechanisms to reason about such threats, whereas PRISM’s multimodal-specific design directly targets this gap.
Appendix I Evaluator-Agnostic Analysis
PRISM employs GPT-4o in two distinct roles: as the SFT data curator that generates chain-of-thought reasoning traces, and as the reward model that scores candidate reasoning steps during MCTS. Crucially, both roles are grounded to ground truth labels from the original dataset: the SFT annotation is conditioned on human-annotated safety violation categories and their associated safety policy specifications, requiring generated rationales and final decisions to be consistent with those labels; similarly, the MCTS helpfulness judgment scores each candidate step against reference answers derived from the original benchmark. This grounding ensures that neither role relies solely on GPT-4o’s internal safety criteria. We investigate whether these roles nonetheless introduce evaluator-specific artifacts through two controlled experiments targeting each role independently.
Curator-Agnostic SFT Data Generation. We examine whether PRISM-CoT’s reasoning quality is contingent on GPT-4o’s particular reasoning style and safety vocabulary by substituting Qwen2.5-VL, a model from an entirely different organization and training lineage, as the SFT data curator. In both cases, the curator is provided with the same ground truth safety violation categories and policy specifications from the original dataset as explicit conditioning; only the model generating the rationale changes. Qwen2.5-VL generates the structured four-step chain-of-thought traces following the same format specification, while the subsequent MCTS-based DPO stage remains unchanged. As shown in Table 6, replacing the curator yields a final model with MIS-hard ASR of $7.59\%$ versus $6.34\%$ under GPT-4o curation, a marginal degradation of $1.25$ percentage points, while MMMU-Pro accuracy drops by only $0.4$ points ($40.9$ vs. $41.3$). The near-identical post-DPO performance across two curators from different model families demonstrates that PRISM’s safety improvements are driven by the structured reasoning framework itself, rather than by any idiosyncrasy in how GPT-4o articulates safety rationales. It also suggests that the DPO stage is robust to variation in SFT initialization quality.
SFT Data Curator MIS-hard ASR $\downarrow$ MMMU-Pro Acc $\uparrow$
GPT-4o 6.34 41.3
Qwen2.5-VL (Self-Curated) 7.59 40.9
Table 6: Replacing the SFT data curator from GPT-4o with Qwen2.5-VL yields near-identical post-DPO performance (MIS-hard ASR difference of 1.251.25pp; MMMU-Pro difference of 0.40.4pts), confirming that improvement is attributable to the structured reasoning framework rather than to curator-specific artifacts.
Cross-Judge Consistency of MCTS Reward Signals. The MCTS preference pairs by construction encode pairwise ordering judgments: one reasoning path is labeled preferred over another. Importantly, each safety judgment is itself grounded to the human-annotated safety violation category and policy specification from the original dataset, and each helpfulness judgment is scored against the reference answer from the original benchmark, rather than relying on GPT-4o’s unconstrained assessment. If the resulting preference orderings were nonetheless idiosyncratic to GPT-4o’s scoring style, a model trained on these pairs would appear aligned only under GPT-4o evaluation. We directly test this by measuring pairwise preference alignment rates, defined as the fraction of pairs on which two judges agree on the ordering, across four diverse judge models: GPT-4o, GPT-4o-mini, Qwen3-VL, and Gemini-2.5. These judges span three independent organizations (OpenAI, Alibaba, Google), different model scales, and architectures trained on distinct data distributions, making systematic agreement across all pairs unlikely unless the preference signal captures shared, objective safety properties.
As shown in Table 7, all pairwise agreement rates exceed $88.5\%$, with the majority of pairs surpassing $92\%$. Notably, cross-organization pairs such as GPT-4o vs. Qwen3-VL ($92.4\%$) and GPT-4o vs. Gemini-2.5 ($94.0\%$) are comparably high to within-organization pairs. This convergent validity across architecturally and organizationally diverse judges provides strong evidence that the MCTS preference pairs encode reward signals that are grounded in objectively recognizable safety violations rather than in the particular annotation style of any single model. Together, both experiments support the conclusion that PRISM’s safety gains are evaluator-agnostic and structurally robust.
Model GPT-4o 4o-mini Qwen3 Gemini
GPT-4o 100.0 97.1 92.4 94.0
GPT-4o-mini 97.1 100.0 92.0 88.5
Qwen3-VL 92.4 92.0 100.0 90.2
Gemini-2.5 94.0 88.5 90.2 100.0
Table 7: Pairwise preference alignment rates (%) across four judge models drawn from three independent organizations. All rates exceed $88.5\%$, including cross-organization pairs, demonstrating that the MCTS reward signals capture objectively recognizable safety properties rather than GPT-4o-specific annotation artifacts.
Appendix J Failure Case Analysis